OpenStudy (anonymous):

2. Solve. 3x – 5 = 7x – 17 (Points : 5) 2 3 5 6

OpenStudy (anonymous):

x=3

OpenStudy (anonymous):

\[3x – 5 = 7x – 17\] \[3x – 7x = 5– 17\] \[-4x = -12\] \[x = 3\]